Assume the following information: 90-day Ghana interest rate = 4% 90-day South African interest rate = 3% 90-day forward rate of South African rand = GHS 0.3500 Spot rate of South African rand = GHS 0.3550 Assume that the Osei Bonsu Co. in Ghana will need 300,000 rand in 90 days to pay for imports from South Africa. It wishes to hedge this payables position. Would it be better off using a forward hedge or a money market hedge? Substantiate your answer with estimated costs for each type of hedge.


1
Expert's answer
2021-09-24T07:09:07-0400

Forward Hedge:

Forward Hedge = Spot Rate (1 + Domestic Interest Rate)/(1 + Foreign Interest Rate)

"forward hedge = 0.3550×\frac{1+ 0.04 }{1+0.03}"

"=0.3550×\frac{1.04}{1.03}"

"=0.3550\u00d71.0097"

Forward hedge rate ="R 0.3584"


Money Market Hedge:

Spot Rate "=GHS0.3550"

"R1=GHS 0.3550"

R300,000 =

"(R300,000 \u00d7 GHS0.3550)\u00f7R1" "=GHS106,500"

Domestic Currency at spot rate = GHS106,500

domestic currency on deposit at prevailing interest rates:

"GHS106,500×4\%=GHS4,260"

At maturity = GHS 4,260 will be

"\frac{4,260×1}{0.3550} =R12,000"


Market Hedge "=\frac{R12,000}{R106,500} =R0.1127"


By using the money market hedge, the company would have effectively locked in a 90-day forward rate of 0.3584.